# Empirical analysis of Binding Precedent efficiency in the Brazilian Supreme Court via Similar Case Retrieval

### v2

@@ Discussion
One sees from LABEL:tab:similar_case_matching that the TF-IDF models, equipped with the classifiers SVM and logistic regression, consistently achieve the best performance across all BPs, measured by the $F_{1}$ score. On the other hand, the deep learning models LSTM and BERT show slightly lower scores. This phenomenon has already been reported in [47], using, as we do Brazilian legal documents (more precisely, data from Supremo em Números). In addition, it has been observed that the TF-IDF-based models outperform certain more modern embeddings (such as Doc2vec, Universal Sentence Encoder, and Longformer). In this context, TF-IDF’s superior performance is attributed to its ability to leverage specific words that allow it to identify the BP. This is supported by the regex search: although being a rather simple way of answering the problem of Similar Case Retrieval, its performance isn’t too bad, even outperforming deep learning models for BP 11.
The case of TF-IDF coupled with random forest is interesting: it achieves perfect precision across all BPs. That is to say, it is highly reliable when it comes to positive predictions. On the other hand, its recall is the lowest in almost every case. Its high precision will be of great importance in Section 4, when we apply the models to Dataset #2. Indeed, we will observe that the models suffer from a serious generalization problem, more precisely, they detect too many documents, which are not related to the topic of the BP. The model TF-IDF+random forest, on the other hand, will mainly detect documents that we have checked as relevant. This comes at the cost of potentially missing documents; we will therefore be using the other TF-IDF models in a complementary way.
It should be noted that the performances of LSTM and BERT may suffer from the fact that their inputs are limited to 512 tokens. One way around this problem, although not considered in our work, is to cut the input documents into several parts, apply the models to each of these parts, and agglomerate the results — a method known as pooling [76]. In addition, we have used the BERT model as is, trained previously on Brazilian legal documents, but not fine-tuned to our specific dataset. The implementation of these two corrections could significantly improve the performance of these models, and we intend to study them in a subsequent work.
@@ 3.4 Explainability
To understand further what the models have learned, we move on to study their most important features. For the TF-IDF models, common measures of importance can be employed and directly computed from the models. For BERT, however, there is no direct way of identifying the most important features, thus we will use the explainability algorithm LIME [48]. Note that, although not presented here, a similar analysis could be made with LSTM.
@@
For the classifiers SVM, logistic regression and random forest, based on the TF-IDF vectorization, common measures of importance of features are respectively the weights of the linear kernel, the coefficients in the decision function, and the standard deviation of impurity decrease in the trees. We compute these quantities via the native functions of scipy. We inspect, for each model, the top features, and gather those common to all models. LABEL:tab:importance_features presents the importance of the words selected this way.
@@
BP Classifier Features
11 algemas uso algemado audiência nulidade
(handcuffs) (use) (handcuffed) (court hearing) (nullity)
SVM 1.0 0.43 0.39 0.33 0.32
logistic 1.0 0.44 0.39 0.37 0.33
forest 1.0 0.49 0.3 0.42 0.25
14 acesso investigação criminal inquérito documentados
(access) (investigation) (criminal) (inquiry) (documented)
SVM 1.0 0.67 0.61 0.56 0.45
logistic 1.0 0.64 0.55 0.56 0.42
forest 1.0 0.45 0.21 0.58 0.64
17 Guarujá precatório juros mora cálculos
(Guarujá) (court order) (interest) (late) (calculations)
SVM 1.0 0.66 0.62 0.5 0.41
logistic 0.81 1.0 0.96 0.75 0.57
forest 0.01 1.0 0.72 0.52 0.14
26 criminológico progressão execuções regime realização
(criminological) (progression) (executions) (regime) (realization)
SVM 1.0 0.82 0.6 0.58 0.49
logistic 1.0 0.86 0.62 0.58 0.45
forest 1.0 0.59 0.38 0.25 0.4
37 isonomia cabe aumentar vencimentos públicos
(isonomy) (apply) (increase) (salaries) (public)
SVM 1.0 0.67 0.61 0.44 0.42
logistic 1.0 0.59 0.56 0.55 0.42
forest 1.0 0.27 0.78 0.46 0.29
The results of LABEL:tab:importance_features are not surprising: the most important words are already contained in the wording of the BPs, or are related grammatically, except two (audieêcia and Guarujá, studied below). We refer the reader to Sections 4.2, 4.3, 4.4, 4.5 and 4.6 for the wordings, as well as a more detailed analysis of the results. In addition, it is remarkable that all three classifiers display similar importance for each of the words. This shows that they have all faithfully learned the statements of the BPs.
The first feature to stand out is audiência (audience/trial), not directly connected to BP 11. However, as explained in Section 4.2, this precedent was born out of the controversy surrounding the use of handcuffs (algemas) in trials. It is therefore consistent that this word has been recognized as important by the models.
Similarly, now in the context of BP 17, the word Guarujá (municipality in the state of São Paulo) is associated with high importance for the models SVM and logistic (1.0 and 0.81) but not for random forest (0.01). As it turns out, this word is only marginally relevant when it comes to identifying the binding precedent: only 9 documents of the training set contain Guarujá, among which 5 cite BP 17. The importance given to this feature is, therefore, a mere consequence of training the models to maximize the $F_{1}$ score, and carries little legal information.
@@ 17 Guarujá precatório juros mora cálculos
This is precisely the underlying mechanism of LIME (Local Interpretable Model-agnostic Explanations) [48], a popular model-agnostic machine learning explainability method. More precisely, to explain the model decision for a specific text sample, LIME perturbs the input of the classifier by randomly removing words from the text and measures the model output probability. Then, after thousands of disturbances, LIME fits a linear regression to predict the model probability from the word presence. From the linear regression coefficients, we can assign an importance score for each word in the input text regarding their contribution to the model decision.
In practice, we use LIME in the following way: for each document in the collection, we calculate with LIME the $N$ most important words (i.e., words which contribute most to the positive class), where $N$ is a hyperparameter, chosen as $N=5$. We then search, among all the documents belonging to a BP, for the most frequent LIME features, and the number of times they appear. The $N$ most important features are represented in LABEL:tab:LIME_features.
Table 4: For each document, LIME detects the most important features. For each BP, we then select the five most frequent features, and indicate the percentage of documents in which LIME detected them.
@@
11 algemas algemado audiência algemados prisão
(handcuffs) (handcuffed) (court hearing) (handcuffed) (prison)
57.0 15.8 5.7 4.7 4.3
14 acesso inquérito criminal defesa investigação
(access) (inquiry) (criminal) (defense) (investigation)
29.7 10.0 8.4 7.1 4.4
17 precatório juros 100 mora precatórios
(court order) (interests) (100) (late) (court orders)
31.1 24.1 12.7 12.0 10.7
26 criminológico progressão regime RAJ execuções
(criminological) (progression) (regime) (Judicial Region) (executions)
33.2 16.4 6.5 5.1 4.4
37 isonomia reajuste vencimentos 2015 equiparação
(isonomy) (readjustment) (salaries) (2015) (equivalent)
18.3 9.2 7.4 4.8 4.5

### v3

@@ Discussion
One sees from Table 2 that the TF-IDF model, equipped with the classifier SVM, consistently achieves the best performance across all BPs, measured by the $F_{1}$ score, with the exception of BP 37, where it swaps places with Longformer. TF-IDF vectorization also performs well with logistic regression, reaching second place (or tied for first) in BPs 11, 14, and 17. In the case of BP 26, LSTM does better, and for BP 37, both LSTM and Longformer have higher scores. On the other hand, the deep learning models LSTM and Longformer show slightly lower scores overall, but they are still competitive. The strong performance of TF-IDF in this task has already been reported in [48], using, as we do, Brazilian legal documents (more precisely, data from Supremo em Números). Namely, it has been observed that the TF-IDF-based models outperform certain more modern embeddings (such as Doc2vec, Universal Sentence Encoder, and Longformer). In this context, TF-IDF’s superior performance is not simply attributed to the fact that deep learning models are limited by the number of tokens — we saw in Section 3.1 that the limitation of 4096 tokens allows almost all documents to be read in their entirety (see Figs. 5(a) and 6(a)) —, but its ability to leverage specific words that allow it to identify the BP. This is supported by the regex search: although being a rather simple way of answering the problem of Case Classification, its performance is not too bad, even outperforming LSTM for BP 11.
The case of TF-IDF coupled with random forest is interesting: it achieves perfect precision across all BPs. That is to say, it is highly reliable when it comes to positive predictions. On the other hand, its recall is the lowest in every case, not including regex. In other words, this model is exposed to the problem of missing potentially relevant documents. This will be observed in Section 4 when applied to Dataset #2, where its usefulness will then be limited.
@@ 3.4 Explainability
To understand further what the models have learned, we move on to study their most important features. For the TF-IDF models, common measures of importance can be directly computed from the models. For Longformer, however, there is no direct way of identifying the most important features, thus we will use the explainability algorithm LIME [49]. We point out that, although not presented here, a similar analysis could be made with LSTM.
@@
For the classifiers SVM, logistic regression, and random forest, based on the TF-IDF vectorization, common measures of importance of features are respectively the weights of the linear kernel, the coefficients in the decision function, and the standard deviation of impurity decrease in the trees. We compute these quantities via the native functions of scipy. We inspect, for each model, the top features, and gather those common to all models. Table 3 presents the importance of the words selected this way.
BP Model Features
11 algemado algemas audiência nulidade uso
(handcuffed) (handcuffs) (court hearing) (nullity) (use)
SVM 0.41 1 0.28 0.29 0.4
logistic 0.38 1 0.3 0.3 0.41
forest 0.39 0.96 0.37 0.27 0.26
14 acesso criminal documentados inquérito investigação
(access) (criminal) (documented) (inquiry) (investigation)
SVM 1 0.52 0.47 0.45 0.59
logistic 1 0.49 0.43 0.49 0.59
forest 1 0.27 0.54 0.65 0.48
17 cálculos juros mora moratórios precatórios
(calculations) (interest) (late) (moratoriums) (court order)
SVM 0.55 0.88 0.68 0.48 1
logistic 0.57 0.98 0.72 0.52 1
forest 0.18 0.9 0.62 0.68 1
26 criminológico exame execuções progressão regime
(criminological) (examination) (executions) (progression) (regime)
SVM 1 0.48 0.58 0.84 0.58
logistic 1 0.5 0.61 0.87 0.58
forest 1 0.57 0.38 0.75 0.32
37 aumentar isonomia servidores simetria vencimentos
(increase) (isonomy) (servants) (symmetry) (salaries)
SVM 0.66 0.95 0.16 0.89 0.46
logistic 0.58 1 0.2 0.67 0.57
forest 1 0.9 0.41 0.19 0.6
@@
The results of Table 3 are not surprising: the most important words are already contained in the wording of the BPs or are related grammatically, except five (audiência, criminal, inquérito, cálculos, and simetria, studied below). We refer the reader to Sections 5.2, 5.3, 5.4, 5.5 and 5.6 for the wordings, as well as a more detailed analysis of the results. It is remarkable that the three classifiers display similar importance for almost each of the words. This shows that they have all faithfully learned the statements of the BPs.
The first feature to stand out is audiência (audience/trial), undirectly connected to BP 11. As we will explain in Section 5.2, this precedent was born out of the controversy surrounding the use of handcuffs (algemas) in trials. It is therefore consistent that this word has been recognized as important by the models.
The second word, criminal, is naturally associated with BP 14, as its field of action is criminal law. In this context, the word is often used in expressions such as persecução criminal (criminal prosecution), instrução criminal (criminal investigation), or responsabilidade criminal (criminal liability). More precisely, 75, 148, and 35 documents in Dataset #1 cite these expressions, respectively. This amounts to 225 documents, 67.5% of which cite BP 14. On the other hand, it is also not surprising to find the word inquérito (inquiry/investigation) in BP 14’s features, since it can be viewed, in this context, as a synonym for procedimento investigatório (investigative procedure), which is at the heart of the binding precedent’s formulation.
For BP 17, a word not present in its wording is detected as important by the models: cálculos (calculations). In Section 5.4, we will present the context of the creation and application of this binding precedent, which arose from a debate about late payment interests. The use of this word is natural in this context, as in the expression “calculation of interests”. For an example of its usage, Justice Marco Aurélio published a clarification of this precedent under Thesis 96888Thesis 96: “Late payment interest applies to the period between the date of the calculations and the date of the requisition or the precatory order” https://portal.stf.jus.br/jurisprudenciaRepercussao/verAndamentoProcesso.asp?incidente=2598262&numeroProcesso=579431&classeProcesso=RE&numeroTema=96, explicitly using the word cálculos. This thesis is reproduced in 79 documents within the dataset.
Finally, the word simetria appears in the expression princípio da simetria (principle of symmetry), a legal concept that was used in conjunction with BP 37 in the context of a controversy regarding the salaries of the Judiciary (magistratura). This is studied in detail in Section 5.6.
In conclusion, these few examples demonstrate that, as was hoped when Dataset #1 was designed, the models are capable of learning not only the statements of the precedents but also some of their context and the themes they cover. Remarkably, these features are highly interpretable: they are words that cover general themes specific to each BP, rather than purely syntactic peculiarities, which would be characteristic of overfitting.
@@
This is precisely the underlying mechanism of LIME (Local Interpretable Model-agnostic Explanations) [49], a popular Machine Learning explainability method. More precisely, to explain the model decision for a specific text sample, LIME perturbs the input of the classifier by randomly removing words from the text and measures the model output probability. Then, after thousands of disturbances, LIME fits a linear regression to predict the model probability from the word presence. From the linear regression coefficients, one assigns an importance score for each word in the input text regarding their contribution to the model decision.
In practice, we use LIME as follows: for each document in the collection, we calculate with LIME the $N$ most important words (i.e., words which contribute most to the positive class), where $N$ is a hyperparameter, chosen as $N=5$. We then search, among all the documents belonging to a BP, for the most frequent features, and the number of times they appear. The $N$ most important features are represented in Table 4.
@@
11 algemas algemado algemados despacho reclamação
(handcuffs) (handcuffed) (handcuffed, plural) (procedural ruling) (complaint)
66.1 22.3 4.6 4.0 1.9
14 acesso autos criminal inquérito defesa
(access) (elements) (criminal) (inquiry) (defense)
26.3 19.9 17.0 16.4 13.4
17 juros precatório moratórios precatórios mora
(interests) (court order) (moratoriums) (court orders) (late)
53.5 41.8 19.2 18.9 17.5
26 criminológico regime SP exame geral
(criminological) (regime) (State of São Paulo) (examination) (general)
44.5 36.2 19.4 15.4 11.3
37 isonomia vencimentos 2015 magistratura aumentar
(isonomy) (salaries) (2015) (Judiciary members) (to increase)
32.7 13.2 9.0 8.6 8.5
Table 4: For each document, LIME detects the most important features. For each BP, we then select the five most frequent features and indicate the percentage of documents in which LIME detected them.